Toby is n years old and Adrian is d years younger than Toby.
Write down a formula for S where S years is the sum of their ages.

Found 2 solutions by mananth, Theo:
Answer by mananth(16946)

Toby is n years old and Adrian is d years younger than Toby.
toby ---n years
adrian's age = n-d
S = n +n-d
S= 2n-d

Answer by Theo(13342)

n = toby's age.
n - d = adrian's age.
the sum of their ages is equal to n + n - d = 2n - d
when the sum of their ages is equal to S, that formula is S = 2n - d
when S = 5, that formula becomes 5 = 2n - d
that should be your answer.

please keep in mind that, if the sum of their ages is 5, then there is a limit for the value of n.

that limit is that n can only be between 2.5 and 5.

when n = 2.5, then 5 = 2n - d becomes 5 = 5 - d which makes d = 0
when n = 5, then 5 = 2n - d becomes 5 = 10 - d which makes d = 5.

if n is less than 2.5, say 2, then S = 2n - d becomes S = 4 - d which makes d = negative which means that adrian is older than toby which can't be.

if n is greater than 5, then the sum of their ages can only be 5 if adrian's age is less than 0 which can't be.

the formula is good, but only if n is between 2.5 and 5.

just keep in mind that there are limits to the value of n for the answer to be valid because toby's age and adrian's age must both be greater than or equal to 0.

those limits are that n can only be between 2.5 and 5.

since they only asked you for the formula, then your answer is that the formula is 5 = 2n - d